OpenStudy (anonymous):

Area under y = arctan(x) between x = 1 and x = 0

OpenStudy (anonymous):

I don't understand how to integrate arctan(x) by hand

OpenStudy (lgbasallote):

i think i remember it having something to do with integration by parts...

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

We have to somehow use the tan function though

OpenStudy (anonymous):

It says use the fact that y=artan(x) is the inverse of y = tan(x) to calculate the area between x = 1 and x =0

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

We can also graph it for a hint

OpenStudy (anonymous):

But we haven't learned integration by parts yet

sam (.sam.):

@baddinlol You have to use integration by parts here, \[\int\limits \tan ^{-1}(x) \, dx\] \[u=\tan^{-1}x~~~~~dv=dx \\ \\ du=\frac{1}{1+x^2} ~~~~~v=x\] \[x \tan ^{-1}x-\int\limits \frac{x}{1+x^2} \, dx\] Then continue...